To attract retailers to its shopping center, the Marketplace Mall will lend money to tenants under formal contracts, provided that they use it to renovate their store space. On November 1, 2017, the company loaned $109,000 to a new tenant on a one-year note with a stated annual interest rate of 9 percent. Interest is to be received by Marketplace Mall on April 30, 2018, and at maturity on October 31, 2018.

Required:

Prepare journal entries that Marketplace Mall would record related to this note on the following dates: (a) November 1, 2017; (b) December 31, 2017 (Marketplace Mall’s fiscal year-end); (c) April 30, 2018; and (d) October 31, 2018. (If no entry is required for a transaction/event, select "No Journal Entry Required" in the first account field.)
Answer #1
Date General Journal Debit Credit
Nov. 1,2017 Notes receivable 109000
     Cash 109000
Dec. 31,2017 Interest receivable 1635 =109000*9%*2/12
     Interest revenue 1635
Apr. 30,2018 Cash 4905
     Interest receivable 1635
     Interest revenue 3270 =109000*9%*4/12
Oct. 31,2018 Cash 4905 =109000*9%*6/12
     Interest revenue 4905
Cash 109000
     Notes receivable 109000